OpenStudy (anonymous):

Choose the Standard Form of the equation y = −1/7x − 3. 1/7x + y = −3 7x + y = −3 x + 7y = −21 −7x + y = −21

OpenStudy (anonymous):

is it x + 7y = −21

OpenStudy (goformit100):

−7x + y = −21

OpenStudy (anonymous):

−7x + y = −21

jhonyy9 (jhonyy9):

x+7y=-21 so than standard form will be 7y=-21 -x y=-21/7 -x/7 y= (-x/7) -3

jhonyy9 (jhonyy9):

x+7y=-21

jhonyy9 (jhonyy9):

y=-1/7x -3 so here 7x or just 7 is denominator ?

jhonyy9 (jhonyy9):

so if just 7 is denominator so than the st

jhonyy9 (jhonyy9):

so than the standard form is x+7y=-21